Rey works in a semiconductor company where he is tasked to manage semiconductor fabrication. He deduced that for every 100 semiconductors, 5 of which are defective. After considering all scenarios, he made the necessary steps to improve the process.


During a site visit, what would be the probability that none of them are defective if the checker selects 3 semiconductors to be tested?


Answer in percentage form and use four decimal places


1
Expert's answer
2021-05-07T12:01:02-0400

Probability of defective semiconductors-

"p=\dfrac{5}{100}=0.05"

then "q=1-p=1-0.05=0.95"


n=3,

Let X be the random variable and denote the number of defective semiconductor


Probability that none of the semiconductor is defective-


"P(X=0)=^nC_0p^0q^{n-0}"

"=^3C_0(0.05)^0(0.95)^3\\ =0.8573"